OpenStudy (anonymous):

Hi guys. So for my calculus textbook I see that all of the answers for the anti derivatives come out as something like 1/2*x^2. I'm just wondering if there is a specific reason as to why they don't just write x^2/2 instead? Is it going to be a good habit as to what I'm going to learn about soon? Thanks!

OpenStudy (amistre64):

its the inverse of the power rule for derivatives ... d/dx x^n = n x^(n-1) to undo it, add 1 to the exponent and divide n x^(n-1+1) ---------- = x^n n-1+1

OpenStudy (anonymous):

Oh wow thanks again amistre64 lol

OpenStudy (amistre64):

yep